OpenStudy (trantom):

I'm wondering whether gravity is really a meters per second per second PER SECOND force because of the fact that when you get closer to an object's center of mass the gravitational pull gets stronger. The problem is that when an apple falls on the earth, the center gets closer but so do the percieved "sides" of the earth (from the apple's point of view). This might mean that it is being pulled sideways as well and it cancels out so the apple's fall really only needs to be explained in m/s/s. In this case, when the angle between the two "sides" of the dominant.... continued below

OpenStudy (trantom):

gravity well producer is above 90 degrees it cancels out but when it is less than 90 it doesn't affect anything and the effect is in m/s/s/s. Even so, when the side gravitational vectors get wider, they also get STRONGER, so it would pull less down, but down HARDER, and because they are balanced in all directions, my original hypothesis would be correct. Gravity is a meters per second per second per second force, even when the angle between the two sides of the dominant gravitational source is more than 90 degrees. Please help D:

OpenStudy (anonymous):

ok what? :P

OpenStudy (anonymous):

first of all.. when u mean gravity i think u mean gravitational acceleration.. which cannot have m/s/s/s.. cause we don't measure it that way!

OpenStudy (trantom):

in the continued part of my question when I said "HARDER" I meant "STRONGER" so... yea thats probably obvious but I'm trying to be as clear as possible

OpenStudy (anonymous):

Forces are measured in \(N=kg\cdot m/s^2\)

OpenStudy (anonymous):

Gravity is a force.

OpenStudy (anonymous):

1/r^2 dependence of gravitational force itself explains why the force gets stronger and stronger as you get closer and closer to the CM

OpenStudy (trantom):

well for earth G is \[9.8\times \left( 4000\div4000+h \right)\]

OpenStudy (anonymous):

so you are basically saying the force should be 1/r^3?

OpenStudy (anonymous):

what is 4000 ? radius of earth is R = 6400 km!

OpenStudy (trantom):

im talking in miles sorry

OpenStudy (trantom):

I can use kilometers sorry about that

OpenStudy (anonymous):

so at any height h \[g(h) = \frac{g.R}{R+h}\]

OpenStudy (anonymous):

ok ok.. no problem!

OpenStudy (trantom):

well what im asking is do the sides of the earth slow down your descent when you fall?

OpenStudy (trantom):

oh gosh I'm so bad at stating my questions...